12. You are purifying a full-length IgG antibody using recombinant expression in E. coli cells. You are using two expression plasmids, one encoding the heavy chain (53 kDa) and the other encoding the light chain (23 kDa). You have added two different affinity tags to the C-termini of each gene (+5 kDa for the heavy chain and +8 kDa for the light chain). (a) What is the molecular weight of your native antibody? (4 pts) (b) Following affinity purification, you run a gel of your eluate in the presence or absence of β-ME (see below). Label each band in the gel (6 pts)-what do you think is going on here? (2 pts) STD +BME - BME 250 150 100 75 50 37 25 15 10
media%2F4ec%2F4ec97195-edcb-4024-ace8-d2
Answer #1

a. Molecular Weight is 178

Heavy Chain (53+5=58) and Light Chain (23+8=31)=89

Heavy Chain (53+5=58) and Light Chain (23+8=31)=89

                                                                Total                      = 178

b.

(b) Following affinity purification, you run a gel of your eluate in the presence or absence of B-ME (see below). Label eachBeta mercaptoethanol involved in the reduction of disulphide bonds present in the immunoglobulin leads to the formation of heavy chains and light chains.

C. It can bind the anion exchange column

(d) You run another gel. Assuming your purification was successful, draw the result you expect in the presence and absence of

e

53 KD look D 2.3 kD AD haiv